OpenStudy (anonymous):

What is the expression in factored form? x2 + 14x + 48 A. (x + 6)(x - 8) B. (x + 8)(x - 6) C. (x - 8)(x - 6) D. (x + 6)(x + 8)

Parth (parthkohli):

\[\textbf{HINT}\]\[x^2 + 14x + 48 = x^2 + 8x + 6x + 48 = \cdots\]